# Supervision, Writing – review and editing

### v1

@@ 3 Methods - Brain Latent Progression (BrLP)
We now introduce the architecture of BrLP, comprising four key components: an LDM, a ControlNet, an auxiliary model, and a LAS block, each described in successive paragraphs. These four components, summarized in Figure 1, collectively address the challenges outlined in the introduction. In particular, the LDM is designed to generate random 3D brain MRIs that conform to specific covariates, while ControlNet aims to specialize these MRI scans to specific anatomical structures of a subject. Additionally, the auxiliary model leverages prior knowledge of disease progression to improve the precision in predicting the volumetric changes of specific brain regions. Finally, the LAS block is used during inference to improve spatiotemporal consistency, as well as to derive a measure of uncertainty for the predictions.
@@ 3.2 ControlNet - conditioning on subject brain MRI
The LDM provides only a limited degree of control over the generated brain MRI via the covariates $c$, and it does not allow for conditioning the model on individual anatomical structures. The purpose of this block is to extend the capabilities of the LDM to encompass this additional control. To achieve this, we use ControlNet (Zhang et al., 2023), (block C in Figure 1) a neural network designed to work in conjunction with the LDM. We conceptualize ControlNet and LDM as a unified network $\epsilon_{\theta,\phi}$, where $\theta$ represents the fixed network’s parameters of the LDM and $\phi$ denotes the trainable network’s parameters of ControlNet. As in the LDM, $\epsilon_{\theta,\phi}$ is still used to predict the noise $\epsilon_{\theta,\phi}(z_{t},t,c,z)$ in the reverse diffusion step $z_{t}\to z_{t-1}$, now incorporating $z=\mathcal{E}(x)$ as a condition to encompass the structure of the target brain $x$ during the generation process. To train ControlNet, we use the latent representations $z^{(A)}$ and $z^{(B)}$ from pairs of brain MRIs of the same patient taken at different ages $A$ < $B$. The covariates $c^{(B)}$ associated with $z^{(B)}$ are known and used as target covariates. Each training iteration involves: i) sampling $t\sim U[1,T]$, ii) performing $t$ forward diffusion steps $z^{(B)}\to z_{t}^{(B)}$, iii) predicting the noise $\epsilon_{\theta,\phi}(z_{t}^{(B)},t,c^{(B)},z^{(A)})$ to revert $z^{(B)}_{t}\to z^{(B)}_{t-1}$, and iv) minimizing the loss $\mathcal{L}_{\epsilon}$ (Eq. 1).
@@ 3.3 Proposed auxiliary model - leveraging disease prior knowledge
AD-related regions shrink or expand over time and at different rates (Pini et al., 2016). Deep-learning-based spatiotemporal models strive to learn these progression rates directly from brain MRIs in a black-box manner, which can be very challenging. To aid this process, we propose incorporating prior knowledge of volumetric changes directly into our pipeline. To do so, we exploit an auxiliary model $f_{\psi}$ (block D in Figure 1) able to predict how the volumes of AD-related regions change over time and provide this information to the LDM via the progression-related covariates $v$. The choice of our auxiliary model is tailored to two scenarios, making BrLP flexible for both cross-sectional and longitudinal data. For subjects with a single scan available at age $A$, we employ a regression model to estimate volumetric changes $\hat{v}^{(B)}=f_{\psi}(c^{(A)})$ at age $B$. For subjects with $n$ past visits accessible at ages $A_{1},\dots,A_{n}$, we predict $\hat{v}^{(B)}=f\psi(c^{(A_{1})},\dots,c^{(A_{n})})$ using Disease Course Mapping (DCM) (Schiratti et al., 2017; Koval et al., 2021), a model specifically designed for disease progression. DCM is intended to provide a more accurate trajectory in alignment with the subject’s history of volumetric changes available. While we employ DCM as a potential solution, any suitable disease progression model can be used in BrLP.
@@ 3.6 Quantifying the uncertainty of the prediction
We then average the components of $\sigma^{(B)}$ into a scalar uncertainty measure, defined as $u^{(B)}$.
@@ 3.7 Implementation settings
The architectures of the autoencoder, the UNet and the ControlNet are taken from the Generative Models library (Pinaya et al., 2023) available in MONAI (Cardoso et al., 2022). The LDM block, which includes the autoencoder $(\mathcal{E},\mathcal{D})$ and the UNet $\epsilon_{\theta}$, follows the same settings proposed in (Pinaya et al., 2022). The autoencoder is fine-tuned from their available pre-trained model [^1] Model available at https://github.com/Project-MONAI/GenerativeModels/tree/main/model-zoo/models/brain_image_synthesis_latent_diffusion_model. The training is performed using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 8. The encoder $\mathcal{E}$ maps the input 3D brain MRI of shape $122\times 146\times 122$ to latent representations of shape $3\times 16\times 20\times 16$. The UNet $\epsilon_{\theta}$ is randomly initialized and trained using the AdamW optimizer with a learning rate of $2.5\times 10^{-5}$ and a batch size of 16. During training, we use $T=1000$ and a scaled linear $\beta$ noise schedule from $\beta_{1}=0.0015$ to $\beta_{T}=0.0205$. The parameters $\phi$ of the ControlNet are randomly initialized and trained using the AdamW optimizer, with a learning rate of $2.5\times 10^{-5}$ and a batch size of 16. The conditioning on the starting age $A$ is achieved by concatenating $A$ with the latent vector $z^{(A)}$ along the channel axis. The LDM’s parameters $\theta$ remain fixed during this training phase. The settings for $T$ and the noise schedule $b_{t}$ are the same used in the LDM. The training is conducted on a GeForce RTX 4090 GPU with 24GB of VRAM. The training steps have required 51 hours for the autoencoder, 12 hours for the UNet, and 12 hours for the ControlNet, respectively. During inference, we use DDIM sampling (Song et al., 2021) with only 25 denoising steps. For the LAS block, we parallelize the generation of multiple $\hat{z}^{(B)}$ on a single GPU, effectively reducing the overhead.
@@ 3.8 MRI preprocessing
Each T1w brain MRI used in our study is pre-processed using: N4 bias-field correction (Tustison et al., 2010), skull stripping (Hoopes et al., 2022), affine registration to the MNI space, intensity normalization (Shinohara et al., 2014) and resampling to 1.5 mm3. The volumes used as progression-related covariates and for our subsequent evaluation are calculated using SynthSeg 2.0 (Billot et al., 2023) and are expressed as percentages of the total brain volume to account for individual differences.
@@ 4 Experiments and results
In this section, we first describe the datasets and evaluation metrics used in our study. We then present an extensive evaluation of BrLP through five distinct experiments: an ablation study examining BrLP’s components and hyperparameters, a comparative analysis against established baseline methods, an investigation of the impact of cognitive status conditioning, an assessment of our proposed uncertainty metric, and an exploration of BrLP’s potential to reduce Type II errors in clinical trials.
@@ 4.1 Internal and external datasets
We collected a large internal dataset of T1w brain MRIs for training and internal evaluation, as well as an external dataset to further validate our findings on out-of-distribution data. For the internal dataset, we combined three publicly available longitudinal datasets: ADNI 1/2/3/GO (1,990 subjects) (Petersen et al., 2010), OASIS-3 (573 subjects) (LaMontagne et al., 2019), and AIBL (242 subjects) (Ellis et al., 2009). This internal dataset comprises 11,730 T1w brain MRI scans from 2,805 subjects, with each subject having at least two MRIs acquired during different visits. The average time interval between the initial and follow-up visits is 4.3 years (SD = 3.1), with a maximum span of 16 years. Age, sex, and cognitive status data were available for all subjects. The average age is $74\pm 7$ years, and 53% of the subjects are male. Based on the final visit, 43.8% of subjects are classified as CN, 25.7% exhibit or develop MCI, and 30.5% exhibit or develop AD. We randomly split the dataset into a training set (80%), a validation set (5%), and a test set (15%) with no overlapping subjects. The validation set is used for early stopping during training. For the external dataset, we used data from the NACC longitudinal study (Beekly et al., 2007), including 2,257 T1w MRIs from 962 subjects. The average time interval between the initial and follow-up visits is 3.8 years (SD = 2.3), with a maximum span of 13 years. The average age is $68\pm 11$ years, and 38% of the subjects are male. Based on the final visit, 72% of subjects are classified as CN, 14.1% exhibit or develop MCI, and 14% exhibit or develop AD. Data from the external dataset is used solely for evaluation.
Datasets statements: (i) Data was collected by the AIBL study group. AIBL study methodology has been reported previously (Ellis et al., 2009) . (ii) Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The original goal of ADNI was to test whether serial MRI, positron emission tomography (PET), other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of MCI and early AD. The current goals include validating biomarkers for clinical trials, improving the generalizability of ADNI data by increasing diversity in the participant cohort, and to provide data concerning the diagnosis and progression of AD to the scientific community. For up-to-date information, see adni.loni.usc.edu.
@@ 4.3 Ablation studies
First, we evaluate how the LAS hyperparameter $m$ influences BrLP’s predictive accuracy. As shown in Figure 2, our analysis across increasing values of $m$ ($m\in[1,2,4,8,16,32,64]$) reveals consistent improvements in terms of both image-based and volumetric metrics. In particular, increasing $m$ from 2 to 64 leads to gradual improvements: MSE decreases by 7%, volumetric errors reduce by 3% on average, and SSIM improves by 0.68%. Paired t-tests ($\alpha=0.05$) confirm that improvements are statistically significant for all metrics except for volumetric errors in the amygdala and CSF. However, this performance gain comes with increased computational demands. As $m$ increases, so does the computation time, as shown in Figure 2. These results highlight the trade-off between accuracy and available computational resources. Unless otherwise specified, we set $m=64$ for all experiments from this point onward. Next, we conduct an ablation study to assess BrLP’s performance with and without the auxiliary model (AUX) and the LAS. In this experiment, we used the first half of each subject’s MRI visits to predict all subsequent MRIs in the latter half (see sequence-aware settings in Section 4.4 for details). The results demonstrate that both the auxiliary model and the spatiotemporal consistency technique independently improve performance. Specifically, the auxiliary model achieves a 16% reduction in volumetric errors, emphasizing the importance of leveraging longitudinal data if available. Meanwhile, the LAS algorithm contributes an additional 4% reduction. Together, these methods yield the best performance, with a combined 21% reduction in volumetric errors.
@@ 4.4 Quantitative and qualitative comparisons with baseline methods
Table 3: Results from the comparison with baseline methods on the external test set. MAE (± SD) in predicted volumes is expressed as a percentage of total brain volume. The result is marked with a star if it is significantly better than all other methods at the 5% significance level (paired t-test with Bonferroni correction).
@@ 4.5 Evaluating the impact of incorrect conditioning on cognitive status in BrLP predictions
Due to the presence of a prediction bias towards healthy aging, we design an experiment to assess whether BrLP’s predictions are correctly influenced by altering the cognitive status of the input subject. Specifically, we conduct quantitative experiments where all AD subjects are conditioned as if they were CN, and we compare these results to those obtained with the correct conditioning. We perform this experiment both in single-image and sequence-aware settings and using both the internal and external test sets. We report the results for this experiment in Table 4. As expected, our findings show that volumetric errors increase when an incorrect cognitive status is provided, especially in the hippocampal region, which is significantly affected in AD subjects. For the hippocampus, the differences are statistically significant in the single-image setting (both in internal and external test sets), but not in the sequence-aware setting. We hypothesize that in sequence-aware settings, the auxiliary model leverages prior visits to approximate the correct progression rate of AD subjects, which mitigates the impact of incorrect conditioning.
@@
4.6 Analysis of the proposed uncertainty measure
@@
In the next two experiments we evaluate the suitability of the uncertainty measure proposed in Section 3.6. First, we examine the relationship between uncertainty and prediction distance, defined as the temporal gap between the age at input MRI and the target age. Intuitively, uncertainty should increase with prediction distance due to the considerable temporal heterogeneity inherent in disease progression. Second, we analyze the relationship between uncertainty and prediction error using image-based metrics. Also in this case, we expect the error to be higher in predictions with higher uncertainty. We use linear mixed-effects models to investigate on these relationships, and we report the fixed effect coefficient $\beta$ along with the p-value $p$.
4.6.1 Uncertainty increases with prediction distance

### v2

@@ 4 Methods - Brain Latent Progression (BrLP)
We now introduce the architecture of BrLP, comprising four key components: an LDM, a ControlNet, an auxiliary model, and a LAS block, each described in successive paragraphs. These four components, summarized in Figure 1, collectively address the challenges outlined in the introduction. In particular, the LDM is designed to generate random 3D brain MRIs that conform to specific covariates, while ControlNet aims to specialize these MRI scans to specific anatomical structures of a subject. Additionally, the auxiliary model leverages prior knowledge of disease progression to improve the precision in predicting the volumetric changes of specific brain regions. Finally, the LAS block is used during inference to improve spatiotemporal consistency, as well as to derive a measure of uncertainty for the predictions both at the global and voxel level.
@@ 4.2 ControlNet - conditioning on subject brain MRI
The LDM provides only a limited degree of control over the generated brain MRI via the covariates $c$, and it does not allow for conditioning the model on individual anatomical structures. The purpose of this block is to extend the capabilities of the LDM to encompass this additional control. To achieve this, we use ControlNet (Zhang et al., 2023), (block C in Figure 1) a neural network designed to work in conjunction with the LDM. We conceptualize ControlNet and LDM as a unified network $\epsilon_{\theta,\phi}$, where $\theta$ represents the fixed network’s parameters of the LDM and $\phi$ denotes the trainable network’s parameters of ControlNet. As in the LDM, $\epsilon_{\theta,\phi}$ is still used to predict the noise $\epsilon_{\theta,\phi}(z_{t},t,c,z)$ in the reverse diffusion step $z_{t}\to z_{t-1}$, now incorporating $z=\mathcal{E}(x)$ as a condition to encompass the structure of the target brain $x$ during the generation process. To train ControlNet, we use the latent representations $z^{(A)}$ and $z^{(B)}$ from pairs of brain MRIs of the same patient taken at different ages $A$ < $B$. The ground-truth covariates $c^{(B)}$ associated with $z^{(B)}$ are known during training and used as target covariates. Each training iteration involves: i) sampling $t\sim U[1,T]$, ii) performing $t$ forward diffusion steps $z^{(B)}\to z_{t}^{(B)}$, iii) predicting the noise $\epsilon_{\theta,\phi}(z_{t}^{(B)},t,c^{(B)},z^{(A)})$ to revert $z^{(B)}_{t}\to z^{(B)}_{t-1}$, and iv) minimizing the loss $\mathcal{L}_{\epsilon}$ (Eq. 1).
@@ 4.3 Proposed auxiliary model - leveraging disease prior knowledge
AD-related regions shrink or expand over time and at different rates (Pini et al., 2016). Deep-learning-based spatiotemporal models strive to learn these progression rates directly from brain MRIs in a black-box manner, which can be very challenging. To aid this process, we propose incorporating prior knowledge of volumetric changes directly into our pipeline. To do so, we exploit an auxiliary model $f_{\psi}$ (block D in Figure 1) able to predict how the volumes of AD-related regions change over time and provide this information to the LDM via the progression-related covariates $v$. The choice of our auxiliary model is tailored to two scenarios, making BrLP flexible for both cross-sectional and longitudinal data. For subjects with a single scan available at age $A$, we employ a regression model to estimate volumetric changes $\hat{v}^{(B)}=f_{\psi}(c^{(A)})$ at age $B$. For subjects with $n$ past visits accessible at ages $A_{1},\dots,A_{n}$, we predict $\hat{v}^{(B)}=f_{\psi}(c^{(A_{1})},\dots,c^{(A_{n})})$ using Disease Course Mapping (DCM) (Schiratti et al., 2017; Koval et al., 2021), a model specifically designed for disease progression. DCM is intended to provide a more accurate trajectory in alignment with the subject’s history of volumetric changes available. While we employ DCM as a potential solution, any suitable disease progression model can be used in BrLP.
@@ 4.6 Quantifying the uncertainty of the prediction
We then average the components of $\sigma^{(B)}$ into a scalar global uncertainty measure, defined as $u^{(B)}$. To assess uncertainty at the voxel level, we decode the $m$ latent predictions into the 3D imaging domain and compute voxel-wise variance to derive an uncertainty map $U^{(B)}$:
$U^{(B)}=\frac{1}{m-1}\sum_{i=1}^{m}\left[\mathcal{D}(z_{i}^{(B)})-\mathcal{D}(\mu^{(B)})\right]^{2}.$ (4)
@@ 4.7 Implementation settings
The architectures of the autoencoder, the UNet, and the ControlNet are taken from the Generative Models library (Pinaya et al., 2023) available in MONAI (Cardoso et al., 2022). The LDM block, which includes the autoencoder $(\mathcal{E},\mathcal{D})$ and the UNet $\epsilon_{\theta}$, follows the same settings proposed in (Pinaya et al., 2022). The autoencoder is fine-tuned from their available pre-trained model [^1] Model available at https://github.com/Project-MONAI/GenerativeModels/tree/main/model-zoo/models/brain_image_synthesis_latent_diffusion_model. The training is performed using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 8. The encoder $\mathcal{E}$ maps the input 3D brain MRI of shape $122\times 146\times 122$ into latent representations of shape $3\times 16\times 20\times 16$. The UNet $\epsilon_{\theta}$ is randomly initialized and trained using the AdamW optimizer with a learning rate of $2.5\times 10^{-5}$ and a batch size of 16. During training, we use $T=1000$ and a scaled linear $\beta$ noise schedule from $\beta_{1}=0.0015$ to $\beta_{T}=0.0205$. The parameters $\phi$ of the ControlNet are randomly initialized and trained using the AdamW optimizer, with a learning rate of $2.5\times 10^{-5}$ and a batch size of 16. The conditioning on the starting age $A$ is achieved by concatenating $A$ with the latent vector $z^{(A)}$ along the channel axis. The LDM’s parameters $\theta$ remain fixed during this training phase. The settings for $T$ and the noise schedule $\beta_{t}$ are the same as those used in the LDM. The training is conducted on a GeForce RTX 4090 GPU with 24GB of VRAM. The training steps have required 51 hours for the autoencoder, 12 hours for the UNet, and 12 hours for the ControlNet, respectively. During inference, we use DDIM sampling (Song et al., 2021) with only 25 denoising steps. In Appendix C, we analyze the impact of using fewer denoising steps on BrLP’s performance. For the LAS block, we parallelize the generation of multiple $\hat{z}^{(B)}$ on a single GPU, effectively reducing the overhead.
@@ 4.8 MRI preprocessing
Each T1w brain MRI used in our study is pre-processed using N4 bias-field correction (Tustison et al., 2010), skull stripping (Hoopes et al., 2022), affine registration to the MNI space, intensity normalization (Shinohara et al., 2014) and resampling to 1.5 mm3. The volumes used as progression-related covariates and for our subsequent evaluation are calculated using SynthSeg 2.0 (Billot et al., 2023) and are expressed as percentages of the total brain volume to account for individual differences.
@@ 5 Experiments and results
In this section, we first describe the datasets and evaluation metrics used in our study. We then present an extensive evaluation of BrLP through five distinct experiments: an ablation study examining BrLP’s components and hyperparameters, a comparative analysis against established baseline methods, an investigation of the impact of cognitive status conditioning, an assessment of our proposed uncertainty metrics at the global and voxel level, and an exploration of BrLP’s potential to reduce Type II errors in clinical trials.
Figure 2: Demographic and diagnostic statistics of the internal and external datasets. Distributions include (A) age at baseline, (B) average time interval between the initial and follow-up visits, (C) sex distribution, and (D) diagnosis (CN, MCI, AD) at final visit.
@@ 5.1 Internal and external datasets
We collected a large internal dataset of T1w brain MRIs for training and internal evaluation, as well as an external dataset to further validate our findings on out-of-distribution data. For the internal dataset, we combined three publicly available longitudinal datasets: ADNI 1/2/3/GO (1,990 subjects) (Petersen et al., 2010), OASIS-3 (573 subjects) (LaMontagne et al., 2019), and AIBL (242 subjects) (Ellis et al., 2009). Dataset statements can be found in Appendix B. This internal dataset comprises 11,730 T1w brain MRI scans from 2,805 subjects, with each subject having at least two MRIs acquired during different visits. We randomly split the dataset into a training set (80%), a validation set (5%), and a test set (15%) with no overlapping subjects. The validation set is used for early stopping during training. For the external dataset, we used data from the NACC longitudinal study (Beekly et al., 2007), including 2,257 T1w MRIs from 962 subjects used solely for evaluation. Figure 2 provides an overview of the demographic and diagnostic statistics for both datasets.
@@ 5.3 Ablation studies
First, we evaluate how the LAS hyperparameter $m$ influences BrLP’s predictive accuracy. As shown in Figure 3, our analysis across increasing values of $m$ ($m\in[1,2,4,8,16,32,64]$) reveals consistent improvements in terms of both image-based and volumetric metrics. In particular, increasing $m$ from 2 to 64 leads to gradual improvements: MSE decreases by 7%, volumetric errors reduce by 3% on average, and SSIM improves by 0.68%. Paired t-tests ($\alpha=0.05$) confirm that improvements are statistically significant for all metrics except for volumetric errors in the amygdala and CSF. However, this performance gain comes with increased computational demands. As $m$ increases, so does the computation time, as shown in Figure 3. Furthermore, since we perform the denoising of the $m$ latents in parallel on a single GPU, memory usage increases linearly with $m$: from 4.6 GB at $m=1$ to 9.6 GB at $m=64$ (excluding the memory required for VAE decoding). These results highlight the trade-off between accuracy and available computational resources. Unless otherwise specified, we set $m=64$ for all experiments from this point onward.
Next, we conduct an ablation study to assess BrLP’s performance with and without the auxiliary model (AUX) and the LAS. In this experiment, we used the first half of each subject’s MRI visits to predict all subsequent MRIs in the latter half (see sequence-aware settings in Section 5.4 for details). The results are presented in Table 1. The term “Base” in the table denotes BrLP without LAS (by setting $m{=}1$) and without AUX (i.e., no conditioning on progression-related covariates). The results show that both AUX and LAS independently improve prediction accuracy. Introducing the auxiliary model reduces volumetric error in conditioned regions by an average of 23%, while its effect on unconditioned regions is smaller, yielding a 10% reduction for CSF and no measurable improvement for the thalamus. These findings highlight the benefit of incorporating longitudinal information, and suggest that modeling all available regions with the auxiliary model could further enhance performance. LAS, on the other hand, contributes an additional average reduction of 4% in volumetric error. When combined, AUX and LAS achieve the best overall performance, with a 21% reduction in volumetric prediction error across all evaluated regions.
@@ 5.4 Quantitative and qualitative comparisons with baseline methods
Table 3: Results from the comparison with baseline methods on the external test set. MAE (± SD) in predicted volumes is expressed as a percentage of total brain volume. The result is marked with a star if it is significantly better than all other methods at the 5% significance level (paired t-test).
@@ 5.5 Evaluating the impact of incorrect conditioning on cognitive status in BrLP predictions
Due to the presence of a prediction bias towards healthy aging, we design an experiment to assess whether BrLP’s predictions are correctly influenced by altering the cognitive status of the input subject. Specifically, we conduct quantitative experiments where all AD subjects are conditioned as if they were CN, and we compare these results to those obtained with the correct conditioning. We perform this experiment both in single-image and sequence-aware settings and using both the internal and external test sets. We report the results for this experiment in Table 4. As expected, our findings show that volumetric errors increase when incorrect cognitive status is provided, especially in the hippocampal region, which is significantly affected in AD subjects. While image-based metrics offer limited insight into the localized effects of conditioning, we observe clear and statistically significant differences in key regions such as the hippocampus in the single-image setting across both test sets. In contrast, these differences are not statistically significant in the sequence-aware setting. This suggests that the auxiliary model in the sequence-aware setting might leverage longitudinal information from prior visits to approximate the correct progression rate of AD subjects, thereby mitigating the impact of incorrect conditioning.
@@
Figure 6: Voxel-level uncertainty evaluated for predictions at different timesteps for a single subject. The first and second rows show the ground truth and predicted MRIs, respectively, at each timestep. The third row presents the uncertainty maps, with lighter colors indicating higher uncertainty. The fourth row displays the voxel-level squared error between the ground truth and predicted MRIs.
5.6 Analysis of the proposed uncertainty measures
@@
In the next three experiments we evaluate the suitability of the uncertainty measures proposed in Section 4.6. First, we examine the relationship between global-level uncertainty and prediction distance, defined as the temporal gap between the age at input MRI and the target age. Intuitively, uncertainty should increase with prediction distance due to the considerable temporal heterogeneity inherent in disease progression. Second, we analyze the relationship between global-level uncertainty and prediction error using image-based metrics. Also in this case, we expect the error to be higher in predictions with higher uncertainty. We use linear mixed-effects models to investigate on these relationships, and we report the fixed effect coefficient $\beta$ along with the p-value $p$. In the third experiment, we evaluate the voxel-level uncertainty map quantitatively by measuring its voxel-wise correlation with prediction error, and qualitatively by visualizing how the map evolves with increasing prediction distance.
5.6.1 Global-level uncertainty increases with prediction distance